# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ 5 Real Data Applications
In this section, we evaluate our proposed methods on four real-world datasets, illustrating their practical benefits in diverse online decision-making tasks.
@@
  • Task 1: Online Candidate Screening. The first task is real-time candidate screening for selecting the candidates who can get into the interview process. We use the recruitment dataset [Kaggle, 2020], which contains $N=45,372$ candidates with 11 attributes including education status, handicapped or not and gender. The target binary variable is whether the candidate passes the job interview.
@@
  • Task 2: High-Risk Diabetes Identification. The second task focuses on health screening, where selecting individuals at high risk of diabetes is critical for early intervention. We use the diabetes health indicators dataset [Kaggle, 2021], consisting of $N=70,692$ samples and 22 covariates including demographic attributes (e.g., sex, age, BMI), lifestyle-related features, and several binary health indicators. The target binary variable is whether an individual has diabetes.
@@
  • Task 3: High-Income Individual Selection. The third task involves using the 1994 Census Bureau dataset [Becker and Kohavi, 1996] to identify a subset of individuals with high incomes (i.e., income >50K>50K) for precision marketing purposes. This dataset contains $N=32,561$ records with 14 attributes including gender, race, marital status, education level, and more.
@@
  • Task 4: Airfoil Noise Detection. The final task involves using the airfoil dataset [Brooks et al., 2014] from the UCI Machine Learning Repository to identify samples with high sound pressure. This dataset contains $N=1,503$ observations with five physical covariates (log frequency, angle of attack, chord length, free-stream velocity, and suction side log displacement thickness). The response variable $Y$ represents the scaled sound pressure, and we aim to test $\mathbb{H}_{0t}:Y_{t}\in[-\infty,c)$, where $c$ is the $(1-\pi_{1})$-quantile of $Y$, with $\pi_{1}=0.4$.
@@
For Tasks 1-3, we randomly sample three subsets from the full dataset: $n_{\text{tr}}=1{,}000$ training, $n_{\text{cal}}=1{,}000$ calibration, and $n_{\text{te}}=1{,}000$ test samples. For Task 4, we set $n_{\text{tr}}=300$, $n_{\text{cal}}=300$, and $n_{\text{te}}=1000$. We compare four optimized procedures Opt-SF, Opt-SFS, Opt-LF and Opt-LFS with two benchmarks SAFFRON and LORD++. At each time step, $K=3$ candidate training algorithms (RF, NN, and SVM) were employed, and the model evaluation criterion is ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with $\rho=0.9$. For our methods, we use the optimized conformity score and for the benchmarks, we randomly choose a score at each time $t$. All results are averaged over 500 replications.
@@
Figure 6 summarizes the results across time $t$ for all four tasks. Table 1 reports the empirical FDR and power at the stopping time $T=1,000$. In Tasks 1-3, all methods control the empirical FDR near or below the nominal level, and, Opt-SF consistently achieves the highest power for all cases as expected. However, in task 4, Opt-SF shows mild FDR inflation, whereas Opt-SFS and Opt-LFS maintain valid control at all time points, demonstrating the necessity of the safe feedback strategies in finite-sample error-rate control guarantees. Notably, SAFFRON and LORD++ deliver substantially lower power than our optimized methods in all tasks, underscoring the strength of the OCTF framework.
In sum, our strategies exploit feedback in three key ways: refining FDP estimation within the GAIF framework, updating online calibration sets adaptively, and optimizing model selection. Empirically, these feedback-aware designs consistently outperform existing GAI methods that ignore feedback, achieving substantially higher power across diverse real-data tasks while preserving rigorous online FDR control.
@@
Table 1: Average $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ for different tasks across four datasets (Candidate, Diabetes, Income, Airfoil). The target FDR level is $\alpha=0.3$. Bold numbers represent the best results.
@@ 6 Concluding Remarks
We study online multiple testing with feedback, aiming to develop reliable machine learning methods for real-time decision-making with rigorous FDR and mFDR control. Our key contribution is GAIF, a feedback-enhanced Generalized Alpha-Investing framework that guarantees online FDR control under the standard assumption of independence between null and all other $p$-values. Building on GAIF, we extend the framework to the conformal setting by constructing independent online conformal $p$-values for null hypotheses, and achieve finite-sample mFDR control via our OCTF procedures. To further enhance performance and address distribution shifts among non-nulls, we propose a feedback-driven model selection criterion.
We highlight several potential directions for future work. First, a current limitation of our framework is its focus on distribution shifts in the alternative data. To address alpha-death and piggybacking under such shifts, one can control weighted FDR using forgetting factors or decaying memory mechanism [Ramdas et al., 2017]. Second, extending our framework to accommodate more general forms of distribution shift is another promising direction. One potential approach is to construct online weighted conformal $p$-values [Prinster et al., 2025] and apply them within the OCTF framework. Finally, to establish guarantees under weaker assumptions, one could relax the definition of valid online FDR control and develop new error-rate notions—analogous to the average-coverage criterion studied in online conformal inference [Gibbs and Candès, 2021, Humbert et al., 2025].

### v2

@@ 6 Real Data Applications
In this section, we evaluate our proposed methods on four real-world datasets to illustrate their practical benefits in diverse online decision-making tasks.
@@
  • Task 1: Online Candidate Screening. The first task is real-time candidate screening for selecting the candidates who can get into the interview process. We use the recruitment dataset [Kaggle, 2020], which contains $N=45,372$ candidates with 11 attributes, including education status, handicapped or not, and gender. The target binary variable is whether the candidate passes the job interview.
@@
  • Task 2: High-Risk Diabetes Identification. The second task focuses on health screening, where selecting individuals at high risk of diabetes is critical for early intervention. We use the diabetes health indicators dataset [Kaggle, 2021], which contains $N=70,692$ samples and 22 covariates, including demographic attributes (e.g., sex, age, BMI), lifestyle-related features, and several binary health indicators. The target binary variable is whether an individual has diabetes.
@@
  • Task 3: High-Income Individual Selection. The third task involves using the 1994 Census Bureau dataset [Becker,Barry and Kohavi,Ronny, 1996] to identify a subset of individuals with high incomes (i.e., income >50K>50K) for precision marketing. This dataset contains $N=32,561$ records with 14 attributes, including gender, race, marital status, education level, and more.
@@
  • Task 4: Airfoil Noise Detection. The final task involves using the airfoil dataset [Brooks,Thomas et al., 2014] from the UCI Machine Learning Repository to identify samples with high sound pressure. This dataset contains $N=1,503$ observations with five physical covariates (log frequency, angle of attack, chord length, free-stream velocity, and suction side log displacement thickness). The response variable $Y$ represents the scaled sound pressure, and we aim to test $\mathbb{H}_{0t}:Y_{t}\in[-\infty,c)$, where $c$ is the $(1-\pi_{1})$-quantile of $Y$, with $\pi_{1}=0.4$.
@@
For Tasks 1-3, we randomly sample three subsets from the full dataset: $n_{\text{tr}}=1{,}000$ training, $n_{\text{cal}}=1{,}000$ calibration, and $n_{\text{te}}=1{,}000$ test samples. For Task 4, we set $n_{\text{tr}}=300$, $n_{\text{cal}}=300$, and $n_{\text{te}}=1{,}000$. We compare four optimized procedures Opt-SF, Opt-SFS, Opt-LF and Opt-LFS with two benchmarks SAFFRON and LORD++. At each time step, $K=3$ candidate training algorithms (RF, NN, and SVM) are employed, and the score selection criterion is ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with $\rho=0.9$. Our methods use the optimized conformity score, while the benchmarks use a randomly selected score at each time $t$. All results are averaged over 500 replications.
@@
Figure 8 summarizes the empirical FDR and power over time $t$, and Table 1 reports the corresponding values at the stopping time $T=1,000$ at the target FDR level $\alpha=0.3$. For Tasks 1-3, all methods control the empirical FDR near or below the nominal level, and Opt‑SF consistently achieves the highest power as expected. However, in task 4, Opt-SF shows mild FDR inflation, whereas the safe procedures Opt-SFS and Opt-LFS maintain valid control throughout. This is consistent with our theory in Section 4.3: the safe variants are more conservative, but they provide the finite-sample guarantees needed in difficult settings. Notably, SAFFRON and LORD++ deliver substantially lower power than our optimized methods in all tasks, underscoring the strength of the OCTF framework.
Overall, the proposed feedback-enhanced methods achieve higher power than feedback-ignoring baselines, while the safe variants maintain reliable error control across all four tasks.
@@
Table 1: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ for different tasks across four datasets (Candidate, Diabetes, Income, Airfoil). The target FDR level is $\alpha=0.3$. Bold numbers represent the best results.
@@ 7 Concluding Remarks
We study online multiple testing with feedback, aiming to develop reliable machine learning methods for real-time decision-making with rigorous FDR and mFDR control. Our key contribution is GAIF and its adaptive variants, a feedback-enhanced Generalized Alpha-Investing framework that guarantees online FDR control under the standard assumption of independence between null and all other $p$-values. Building on GAIF, we further develop OCTF for online conformal testing by constructing independent online conformal $p$-values for null hypotheses, which achieves finite-sample mFDR control. To further enhance performance and address distribution shifts among non-nulls, we propose a feedback-driven score selection strategy and provide theoretical analysis characterizing its optimality properties.
We highlight several potential directions for future work. First, the current framework mainly focuses on distribution shifts in the non-null data. To address alpha-death and piggybacking under such shifts, one can consider controlling weighted FDR using forgetting factors or decaying memory mechanism [Ramdas et al., 2017]. Second, extending our framework to accommodate more general forms of distribution shift is another promising direction. One potential approach is to incorporate adaptive strategies that exploit local information, in the spirit of the adaptive conformal inference [Gibbs and Candès, 2021]. Finally, to establish guarantees under weaker assumptions, one could relax the definition of valid online FDR control and develop new error-rate notions, analogous to the average-coverage criterion studied in online conformal inference [Gibbs and Candès, 2021, Humbert et al., 2025].
Acknowledgments
We would like to acknowledge the action editor and three anonymous referees for their valuable comments and suggestions, which have improved the manuscript greatly.